OpenStudy (anonymous):

Which values of x and y satisfy 2x − y = 7 and 3x − 2y = 5 simultaneously?

OpenStudy (kc_kennylau):

"Which" means that you are given choices. Plug in the choices one by one to see if they satisfy the both equations :)

OpenStudy (radar):

This system of equations appears ideal for solving using the "substitution" method. You can do this by taking the first equation 2x - y = 7 and solve for y, you should get y= 2x-7. Now substitute this for y in the second equation (3x-2y=5) after substitution it would look like this: 3x - 2(2x-7) = 5 which becomes 3x -4x + 14 = 5 which becomes -x =-9 or x = 9. Now plug that value for x into the second equation and solve for y.